OpenStudy (bakonloverk):

what would 0.495 be as a fraction please reduce

OpenStudy (anonymous):

okay so first divide by 1

OpenStudy (linknissan):

495/1000

OpenStudy (anonymous):

\[\frac{0.495}{1}\] next multiply the top and bottom by 1000

OpenStudy (bakonloverk):

495 i think

OpenStudy (anonymous):

\[\frac{495}{1000}\] and now you need to reduce

OpenStudy (anonymous):

yeah exactly, 495 for the top, you have to multiply the bottom too though to keep it even

OpenStudy (bakonloverk):

oh ok so then it would be 495 over 1000

OpenStudy (linknissan):

yea...reduce it by dividing 5 in num n ddenomi

OpenStudy (bakonloverk):

so then its 99/200?

OpenStudy (anonymous):

exactly

OpenStudy (bakonloverk):

u can't reduce farther?

OpenStudy (linknissan):

yea rights thts it

OpenStudy (anonymous):

nope. think about it.if instead of 0.495 you had 0.500 you would get the fraction 100/200 or 1/2. so the difference your dealing with is the 0.005. which is 1/200. 0.500 - 0.005 = 0.495. 100/200 - 1/200 = 99/200

OpenStudy (anonymous):

that may just confuse...

OpenStudy (bakonloverk):

oh oko thank you so much u really helped me

OpenStudy (bakonloverk):

no i didnt confuse me i understood it very well thank you

OpenStudy (anonymous):

great :)